# bibliography.bib

### v1

@@ 4.1 Lower bound on the Rademacher Complexity Analysis
A natural question that might come up is whether the obtained bound is tight or not. In this section, we provide an answer to this by showing that the Rademacher complexity is lower bounded in a similar fashion for a class of networks.
@@
Consider the hypothesis space ${\mathcal{H}}$ defined in 4. If $P(\cdot)$ is the average pooling operation, $\sigma(\cdot)$ is the ReLU activation function, then there is a data distribution such that the Rademacher complexity is lower bounded by $c\frac{b_{x}M_{1}M_{2}}{\sqrt{m}}$ where $c$ is an independent constant.
The proof is provided in the Appendix E. Note that Rademacher complexity (RC) bounds are known to be tight for shallow models, for example, for SVMs. Indeed, if one fixes the first layer and only trains the last linear layer with weight decay, the model is equivalent to hard-margin SVM, for which RC bounds are tight. We also provide additional evidence in the numerical result section.
@@ 6 Generalization Bounds for Weight Sharing
In this section, we try to answer the question whether the gain of our bound lies in the specific equivariant architecture or weight sharing. The answer is subtle. Indeed, an arbitrary type of weight sharing will not bring out the generalization gain. However, there are weight-sharing schemes that can lead to a similar gain without necessarily being equivariant.
@@
The following proposition provides a purely norm based generalization bound.
@@
For the class of functions $h$ in the hypothesis space ${\mathcal{H}}^{w.s.}$ with the average pooling operation $P$, $\sigma(\cdot)$ as a $1$-Lipschitz positive homogeneous activation function, the generalization error is bounded as:
@@
Note that if ${\bm{b}}_{k,l}$’s are the rows of the circulant matrices, then we end up with the same result as before with $M^{w.s.}_{2}=M_{2}$. Interestingly, if the vectors $\{{\bm{b}}_{k,l}:k\in[|G|]\}$’s are orthogonal to each other for each $l\in[|G|]$, and all of them have unit norm, i.e., $\left\|{\bm{b}}_{k,l}\right\|=1$, then we also get $M^{w.s.}_{2}=M_{2}.$ The conclusion is that the weight sharing can impact the generalization similarly to the group convolution, even if the construction does not arise from the group convolution. This does not generally hold, particularly if the row vectors of matrices ${\bm{B}}_{k}$ are not orthogonal. On the other hand, the gain of equivariance shows itself for distributions with built-in symmetries.
@@ 7 Generalization Bounds for Local Filters
Consider the hypothesis space of functions ${\mathcal{H}}^{\Phi}$ constructed by patches $\Phi$, the average pooling operation $P$, $1$-Lipschitz positive homogeneous activation function $\sigma$. Let $O_{\Phi}$ be the maximal number of patches that any input entry appears in. Then, with probability at least $1-\delta$ and for all $h\in{\mathcal{H}}^{\Phi}$, we have:
@@
Consider the hypothesis space of group convolution networks with band-limited filters. If the filters have $B$ non-zero entries in the frequency domain, then the generalization error term in Proposition eq. 7.1 for the smallest spatial filters is at least of order ${\mathcal{O}}(b_{x}M_{1}M_{2}/\sqrt{mB})$.
@@ 8 Numerical Results
Let us first introduce the two family of groups we consider in our experiments: the cyclic group $C_{n}$ containing $n$ discrete rotations and the dihedral group $D_{n}$ of $n$ discrete rotations and $n$ reflections.
@@
In this section, we validate our theoretical results on a simpler variation of the rotated MNIST dataset where we consider the binary classification task of determining whether a digit is smaller or larger than $4$. Rather than working with raw images, we pre-process the dataset by linearly projecting each MNIST image via a fixed randomly initialized $G=D_{32}$-steerable convolution layer [Weiler2019] to a single $6400$-dimensional feature vector - interpreted as a $c_{0}=100$-channels signal over $G=D_{32}$; see Appendix K for more details. When constructing equivariance to a subgroup $G<D_{32}$, this feature vector is interpreted as a $c_{0}=100\cdot\frac{|D_{32}|}{|G|}$-channels signal over $G$. We design our equivariant networks as in Eq. 3 and use $c_{1}\cdot|G|\approx 2000$ total intermediate channels for all groups $G$. The dataset is then augmented with random $D_{32}$ transformations to make it symmetric. Hence, we expect that $G$-equivariant models with $G$ closer to $D_{32}$ to perform best.
@@
In these experiments, we are interested in how well our theoretical bound correlates with the observed trends. In particular, we focus on the $\frac{M_{1}M_{2}}{\sqrt{m}}$ term which is the main component in our bounds. In Fig 2(a) and 2(b), we evaluate the networks on different values of training set size $m$ and on different equivariance groups $G$. We first observe that the $\frac{M_{1}M_{2}}{\sqrt{m}}$ term correlates with the empirical generalization error in Fig. 2(a). In particular, larger groups $G$ achieve both lower generalization error and lower value of our norm bound. We also note that this term decreases as $1/\sqrt{m}$ in Fig. 2(b), which indicates that our bound becomes non-vacuous as $m$ increases, and we do not suffer from deficiencies of other norm based bounds [nagarajan_uniform_2019]. Overall, these results suggest that our bound can explain generalization across different choices of equivariance group $G$ as well as different training settings.
(a) The generalization error versus the bound
(b) The bound versus the number of samples
Figure 2: Numerical results for the generalization error
@@ 9 Conclusion
In this work, we provided various generalization bounds on equivariant networks with a single hidden layer. We have used Rademacher complexity analysis to derive these bounds. They are all dimension-free, as they do not depend on the input and output dimensions. They only depend on the norm of learnable weights. We have considered equivariant models in spatial and in frequency domain, models with weight sharing and local filters. The first insight from our analysis is that the suitable weight-sharing techniques should be able to provide similar guarantees. This is not surprising, as we did not assume any symmetry in the data distribution. Other works in the literature have highlighted the benefit of equivariance if such symmetries exist - see for example [sannai_improved_2021, sokolic_generalization_2017]. Finally, local filters can potentially provide an additional gain, although the story for band-limited filters is more subtle. We also provide a lower bound on Rademacher complexity and showed numerically that our generalization error correlates well with the true error and decreases with $m$. In future works, it is interesting to generalize the current result to deeper networks. Besides, we have focused on positively homogeneous activation functions. We expect that the result can be extended to general Lipschitz activations using techniques in [vardi_sample_2022]. However, the current proof techniques would not work for norm-based nonlinearities used in equivariant literature [Worrall2017-HNET, 3d_steerableCNNs, Weiler2019]. We encourage readers to consult Appendix J for comparison with other works and future directions.
@@ Appendix A An Overview of Representation Theory and Equivariant Networks
The order of a group $G$ is the cardinality of its set and is indicated by $|G|$. A group $G$ is finite when $|G|\in{\mathbb{N}}$, i.e., when it has a finite number of elements. A compact group is a group that is also a compact topological space with continuous group operation. Every finite groups is also compact (with the discrete topology).
@@
For example, a group can act on functions overs the group’s elements: given a signal $x:G\to\mathbb{R}$, the action of $g\in G$ on $x$ is defined as $[g.x](h):=x(g^{-1}h)$, i.e. $g$translates" the function $x$.
@@
The most simple representation is the trivial representation $\psi:G\to\mathbb{R},g\mapsto 1$, mapping all element to the multiplicative identity $1\in V=\mathbb{R}$. The common $2$-dimensional rotation matrices are an example of representation on $V=\mathbb{R}^{2}$ of the group $\operatorname{SO}(2)$:
@@
If $V=\mathbb{C}^{|G|}$, then $x,w:G\to\mathbb{C}$ can be interpreted as signals over the group and group convolution takes the more familiar form
@@ A.1 Equivariant Networks
In this work, we consider real valued networks with $2$ layers, i.e. we limit our discussion to an hypothesis class of the form
@@
The representation theoretic tools we introduced earlier allow for a quite general framework to describe a wide family of equivariant networks beyond group-convolution, by considering the generalization of the Fourier transform we discussed in the previous section. By using Theorem A.5, a representation $\rho_{l}$ (for $l=0,1$) decomposes into direct sum of irreps as follows:
@@
This induces a block structure in $\hat{{\bm{W}}}_{l}$, where the $(\psi,i;\psi^{\prime},j)$-th block block maps the $i$-th input block transforming according to $\psi$ to the $j$-th output block transforming under $\psi^{\prime}$, with $i\in[m_{l-1,\psi}]$ and $j\in[m_{l,\psi^{\prime}}]$. By Theorem A.4, there are no linear maps equivariant to $\psi$ and $\psi^{\prime}$ whenever these representations are in-equivalent; this implies the matrix $\hat{{\bm{W}}}_{l}$ is sparse, since its $(\psi,i;\psi^{\prime},j)$-th block is non-zero only when $\psi\cong\psi^{\prime}$. Then, the non-zero $(\psi,i,j)$-th block (here, we drop the redundant $\psi^{\prime}$ index) is denoted by $\hat{w}_{i,j}(\psi)$ and should commute with $\psi(g)$ for any $g\in G$. This is achieved by expressing this matrix as a linear combination of a few fixed basis matrices spanning the space; see [behboodi_pac-bayesian_2022, cesa2022a] for more details.
@@
Each module in an equivariant network should commute with the action of the group on its on input in order to guarantee the overall equivariance of the model. Hence, the activation function $\sigma$ used in the intermediate layer should be equivariant with respect to $\rho_{1}(g)$ as well, i.e. $\sigma(\rho_{1}(g){\bm{x}})=\rho_{2}(g)\sigma({\bm{x}}),\forall g\in G$. There are different activation functions used in the literature. As we use unitary representations, one can use norm nonlinearities, $\sigma({\bm{x}})=\eta(\left\|{\bm{x}}\right\|)\frac{{\bm{x}}}{\left\|{\bm{x}}\right\|}$ for a suitable function $\eta:\mathbb{R}^{+}_{0}\to\mathbb{R}^{+}_{0}$. It can be seen that $\sigma(\rho_{1}(g){\bm{x}})=\rho_{1}(g)\sigma({\bm{x}})$. Some examples are norm ReLU [Worrall2017-HNET], squashing [sabour_dynamic_2017] and gated nonlinearities [Weiler2018-STEERABLE].
@@
The last layer maps the features after the activation to a single scalar and is assumed to be invariant. This linear layer can be thought as a special case of the general framework above with the choice of output representation $\rho_{2}$ being a collection of trivial representations, i.e. $\rho_{2}(g)=I$ is the identity. Since the trivial representation is an irrep, by Theorem A.4, this linear map only captures the invariant information in the output of the activation function. In the case of a GCNN with pointwise activation $\sigma$, this layer acts as an averaging pooling operator over the $|G|$ entries of each channel and learns the weights to linearly combines the $c_{1}$ independent channels. Note that, one can also include a custom pooling operator $P(\cdot)$ (e.g. max-pooling) between the activation layer $\sigma$ and the final layer ${\bm{u}}$ like in eq. 3: this gives additional freedom to construct the final invariant features beyond just average pooling.

### v2

@@ 4.2 Lower bound on the Rademacher Complexity Analysis
A natural question that might come up is whether the obtained bound is tight or not. In this section, we provide an answer to this by showing that the Rademacher complexity is lower bounded similarly for a class of networks.
@@
Consider the hypothesis space ${\mathcal{H}}$ defined in eq. 4. If $P(\cdot)$ is the average pooling operation, and $\sigma(\cdot)$ is the ReLU activation function, then there is a data distribution such that the Rademacher complexity is lower bounded by $c\frac{b_{x}M_{1}M_{2}}{\sqrt{m}}$ where $c$ is an independent constant.
The proof is provided in the Appendix E. Note that Rademacher complexity (RC) bounds are known to be tight for shallow models, such as SVMs. Indeed, if one fixes the first layer and only trains the last linear layer with weight decay, the model is equivalent to hard-margin SVM, for which RC bounds are tight. We also provide additional evidence in the numerical result section.
@@ 6 Generalization Bounds for Weight Sharing
In this section, we try to answer the question whether the gain of our bound lies in the specific equivariant architecture or weight sharing. The answer is subtle. Indeed, an arbitrary type of weight sharing will not bring out the generalization gain. However, weight-sharing schemes can lead to a similar gain without necessarily being equivariant.
@@
The following proposition provides a purely norm-based generalization bound.
@@
For the class of functions $h$ in the hypothesis space ${\mathcal{H}}^{w.s.}$ with the average pooling operation $P$, $\sigma(\cdot)$ as a $1$-Lipschitz positively homogeneous activation function, the generalization error is bounded as:
@@
Note that if ${\bm{b}}_{k,l}$’s are the rows of the circulant matrices, then we end up with the same result as before with $M^{w.s.}_{2}=M_{2}$. Interestingly, if the vectors $\{{\bm{b}}_{k,l}:k\in[|G|]\}$’s are orthogonal to each other for each $l\in[|G|]$, and all of them have unit norm, i.e., $\left\|{\bm{b}}_{k,l}\right\|=1$, then we also get $M^{w.s.}_{2}=M_{2}.$ The conclusion is that the weight sharing can impact the generalization similarly to the group convolution, even if the construction does not arise from the group convolution. This does not generally hold, particularly if the row vectors of matrices ${\bm{B}}_{k}$ are not orthogonal. Note that throughout the paper, we have not assumed anything regarding the underlying distribution, which can be without symmetry. On the other hand, the gain of equivariance shows itself for distributions with built-in symmetries if we assume that the underlying task has invariance or equivariance property.
@@ 7 Generalization Bounds for Local Filters
Consider the hypothesis space of functions ${\mathcal{H}}^{\Phi}$ constructed by patches $\Phi$, the average pooling operation $P$, $1$-Lipschitz positively homogeneous activation function $\sigma$. Let $O_{\Phi}$ be the maximal number of patches that any input entry appears in. Then, with probability at least $1-\delta$ and for all $h\in{\mathcal{H}}^{\Phi}$, we have:
@@
Consider the hypothesis space of group convolution networks with band-limited filters. If the filters have $B$ non-zero entries in the frequency domain, then the generalization error term in Proposition 7.1 for the smallest spatial filters is at least of order ${\mathcal{O}}(b_{x}M_{1}M_{2}/\sqrt{mB})$.
@@
(a) Generalization error vs our bound on MNIST
(b) Our bound vs number of samples on MNIST
(c) Generalization error vs our bound on CIFAR10
(d) Our bound vs number of samples on CIFAR10
Figure 2: Numerical results for the generalization error on the rotated MNIST and CIFAR10 datasets. The plots on the left, (a)-(c), confirm that our theoretical bound captures the effect of different configurations - equivariance groups $G$, training set sizes $m$ and datasets - on the generalization error, i.e. there is a positive correlation between the generalization error and our theoretical bound $\frac{M_{1}M_{2}}{\sqrt{m}}$ across all these cases. On the right, (b)-(d), we verify the bound decreases following a trend similar to $\frac{1}{\sqrt{m}}$ and approaching zero for large training set sizes $m$.
@@ 8 Numerical Results
Let us first introduce the two family of groups we consider in our experiments: the cyclic group $C_{n}$ containing $n$ discrete rotations and the dihedral group $D_{n}$ of $n$ discrete rotations and $n$ reflections (see Def. A.2 and Def. A.3 for precise definitions of these groups).
@@
In this section, we validate our theoretical results on a variation of the rotated MNIST and CIFAR10 datasets where we consider a simpler binary classification task (determining whether a digit is smaller or larger than $4$ in MNIST and whether an image belongs to the first $5$ classes of CIFAR10 or not). Rather than working with raw images, we pre-process the dataset by linearly projecting each image via a fixed randomly initialized $G=D_{32}$-steerable convolution layer [Weiler2019] to a single $6400$-dimensional feature vector333 This is a random linear transformation of the input images into feature vectors: the reader can think of this as analogous to reshaping an image before feeding it into an MLP, but with some care to preserve rotation and reflection equivariance. Note also that, since the output ($6400$) is much larger than the number of input pixels, this transformation is practically invertible and preserves all relevant input information. - interpreted as a $c_{0}=100$-channels signal over $G=D_{32}$; see Appendix K for more details and Fig. 5 for a visualization of this linear projection. When constructing equivariance to a subgroup $G<D_{32}$, this feature vector is interpreted as a $c_{0}=100\cdot\frac{|D_{32}|}{|G|}$-channels signal over $G$. We design our equivariant networks as in Eq. 3 and use $c_{1}\cdot|G|\approx 2000$ total intermediate channels for all groups $G$ in the MNIST experiments and $c_{1}\cdot|G|\approx 4000$ in the CIFAR10 ones. The dataset is then augmented with random $D_{32}$ transformations to make it symmetric. Hence, we expect that $G$-equivariant models with $G$ closer to $D_{32}$ to perform best.
@@
In these experiments, we are interested in how well our theoretical bound correlates with the observed trends. In particular, we focus on the $\frac{M_{1}M_{2}}{\sqrt{m}}$ term which is the main component in our bounds. In Fig 2, we evaluate the networks on different values of training set size $m$ and on different equivariance groups $G$. We first observe that the $\frac{M_{1}M_{2}}{\sqrt{m}}$ term correlates well with the empirical generalization error on both the MNIST dataset in Fig. 2(a) and the CIFAR10 dataset in Fig. 2(c). In particular, larger groups $G$ achieve both lower generalization error and lower value of our norm bound. We also note that this term decreases as $1/\sqrt{m}$ in Fig. 2(b), which indicates that our bound becomes non-vacuous as $m$ increases, and we do not suffer from deficiencies of other norm based bounds [nagarajan_uniform_2019].
Overall, these results suggest that our bound can explain generalization across different choices of equivariance group $G$, as well as different training settings.
Next, we investigate how other equivariant design choices affect generalization and verify if our bound can model this effect. Precisely, when parameterizing the filters over the group $G$ in the frequency (i.e. irreps) domain (as in Sec 5), it is common to only use a smaller subset of all frequencies. A bandlimited sub-set of frequencies can be interpreted as a form of locality in the frequency domain, in analogy to the local filters in Sec 7. In Fig.3, we experiments on CIFAR10 with different variations of our equivariant architecture obtained by varying the maximum frequency used to parameterize filters. In these experiments, we keep the training dataset size fixed to $m=3200$ and we only consider the largest groups (since a wider range of frequencies in $[0,\dots,N/2]$ can be chosen).
In Fig.3, we find that our bound captures the effect of locality in the frequency domain on the generalization.
Figure 3: Our bound vs. the generalization error on CIFAR10 (training set size $m=3200$) when varying the maximum frequency used to parameterize the filters, which is the locality effect in the frequency domain. Each dot and its error bars represent the mean and standard deviation over at least $3$ runs with the same configuration. As expected, architectures leveraging a lower frequency design achieve lower generalization error and our bound can exactly capture this effect. Note that increased frequencies are still beneficial for the final test performance: we study the trade-off between generalization and test performance as a function of the model frequency in Fig. 4.
In Fig. 3, we observed that higher frequencies typically lead to worse generalization. For this reason, we include a final study to explore the effect of locality in the frequency domain - i.e., band-limitation - on the final test performance of the models and the associated trade-offs between generalization and expressivity. Fig. 4 compares the generalization error with the final test accuracy of some of our models when varying the maximum frequency of the filters and the training set size $m$. This visualization highlights a particular behavior: low-frequency models tend to achieve lower generalization error because of worse data fitting (lower test accuracy). Conversely, higher frequency models often show higher generalization errors (especially in the low data regime) but achieve much higher test performance. Moreover, high-frequency models benefit the most from increased dataset size $m$. These properties lead to the (multiple) V-shaped patterns in Fig. 4: on the smaller datasets (light dots in Fig. 4), while increasing frequencies leads to improved test accuracy, it also increases the generalization error. On the other hand, in the larger dataset (dark dots in Fig. 4), higher frequencies directly improve the test performance and even show minor improvements in generalization error. When varying dataset size, these different behaviors form multiple V-shaped patterns in Fig. 4 (highlighted with the dotted lines).
Figure 4: Test accuracy vs. Generalization error on CIFAR10 when varying the maximum frequency used to parameterize the filters and the training set size $m$. Each dot is the average performance over at least $3$ runs with the same configuration. In Fig. 3, we found that higher frequencies correlated with higher generalization error: here, we study the trade-off between generalization and test performance. For each dataset size $m$, increasing the frequency improves the test performance until a certain saturation point; beyond that, increased frequencies mostly lead to increased generalization error. We highlight this effect by drawing four dotted lines following the trends for varying frequencies with $G=C_{32}$ on four different dataset sizes $m$: the different slopes of the curves correspond to the different saturation effects.
8.1 Main Insights from Numerical Results
We would like to summarize some of the main insights from the numerical examples, some of which pose interesting theoretical questions.
The generalization error seems to be lowest at the medium group sizes as seen in MNIST or CIFAR experiments, although the trend is unclear. There seems to be a group size that leads to the lowest generalization error, and if this conjecture is correct is an open question.
The other observation is about the relation between test accuracy and the generalization error. In general, it can be seen that the highest test accuracies have lower generalization errors in all the plots in Figure 4. It is interesting to directly explore the impact of the group size on the test error, which requires a different machinery.
@@ 9 Conclusion
In this work, we provided various generalization bounds on equivariant networks with a single hidden layer as well as a bound for multi-layer neural networks. We have used Rademacher complexity analysis to derive these bounds, where the proofs were based on either direct analysis of the Rademacher complexity or covering number arguments. The bounds are mostly dimension-free, as they do not depend on the input and output dimensions. They only depend on the norm of learnable weights. The situation changes for multi-layer scenarios and max-pooling where some dimension dependency appears in the bound. In light of the results on the limitation of uniform complexity bounds for deeper neural networks [nagarajan_uniform_2019, jiang_fantastic_2020], we believe that the bound on multi-layer scenario would suffer from similar limitations. However, the emergence of some dimensions in the bound for max-pooling seems related to how covering number argument was applied. Whether the dependency can be removed for max-pooling is an interesting research direction.
We have considered equivariant models in spatial and frequency domains, models with weight sharing, and local filters. The first insight from our analysis is that suitable weight-sharing techniques should be able to provide similar guarantees. This is not surprising, as we did not assume any symmetry in the data distribution. Other works in the literature have highlighted the benefit of equivariance if such symmetries exist - see, for example, [sannai_improved_2021, sokolic_generalization_2017]. Finally, local filters can potentially provide an additional gain, although the story for band-limited filters is more subtle. We also provide a lower bound on Rademacher complexity. We have conducted extensive numerical experiments and investigated the correlation between our generalization error and the true error, as well as the relation between the number of samples, group size, and frequency.
We have focused on positively homogeneous activation functions. We expect that the result can be extended to general Lipschitz activations using techniques in [vardi_sample_2022]. However, the current proof techniques would not work for norm-based nonlinearities used in equivariant literature [Worrall2017-HNET, 3d_steerableCNNs, Weiler2019]. We encourage readers to consult Appendix J for comparison with other works and future directions.
@@ Appendix A An Overview of Representation Theory and Equivariant Networks
The order of a group $G$ is the cardinality of its set and is indicated by $|G|$. A group $G$ is finite when $|G|\in{\mathbb{N}}$, i.e. when it has a finite number of elements. A compact group is a group that is also a compact topological space with continuous group operation. Every finite group is also compact (with a discrete topology).
We now present two examples of finite groups that we used throughout our experiments, the cyclic and the dihedral groups.
Definition A.2 (Cyclic Group).
The cyclic group $C_{N}$ of order $N\in{\mathbb{N}}$ is the group of $N$ discrete rotations by angles which are integer multiples of $\frac{2\pi}{N}$, i.e. $\{R_{p\frac{2\pi}{N}}\ |\ p\in[0,1,\dots,N-1]\}$.
The binary operation combines two rotations to generate the sum of the two rotations. This is represented by integer sum modulo $N$, i.e.
$R_{p\frac{2\pi}{N}}\cdot R_{q\frac{2\pi}{N}}=R_{(p+q\mod N)\frac{2\pi}{N}}$
Definition A.3 (Dihedral Group).
The dihedral group $D_{N}$ of order $2N\in{\mathbb{N}}$ is the group of $N$ discrete rotations (by angles which are integer multiples of $\frac{2\pi}{N}$) and $N$ reflections (generated by a reflection along an axis followed by any of the $N$ rotations), i.e.
$\{R_{p\frac{2\pi}{N}}\ |\ p\in[0,1,\dots,N-1]\}\cup\{R_{p\frac{2\pi}{N}}F\ |\ p\in[0,1,\dots,N-1]\}$
where $F$ is a reflection along an axis. Note that the group $D_{N}$ has size $2N$ and contains the group $C_{N}$ as a subgroup.
Another important concept is that of group action:
@@
For example, a group can act on functions over the group’s elements: given a signal $x:G\to\mathbb{R}$, the action of $g\in G$ on $x$ is defined as $[g.x](h):=x(g^{-1}h)$, i.e., $g$translates" the function $x$.
@@
The most simple representation is the trivial representation $\psi:G\to\mathbb{R},g\mapsto 1$, mapping all elements to the multiplicative identity $1\in V=\mathbb{R}$. The common $2$-dimensional rotation matrices are an example of representation on $V=\mathbb{R}^{2}$ of the group $\operatorname{SO}(2)$:
@@
If $V=\mathbb{C}^{|G|}$, then $x,w:G\to\mathbb{C}$ can be interpreted as signals over the group, and group convolution take the more familiar form
@@ A.1 Equivariant Networks
In this work, we consider real valued networks with $2$ layers, i.e., we limit our discussion to a hypothesis class of the form
@@
The representation theoretic tools we introduced earlier allow for a quite general framework to describe a wide family of equivariant networks beyond group-convolution by considering the generalization of the Fourier transform we discussed in the previous section. By using Theorem A.7, a representation $\rho_{l}$ (for $l=0,1$) decomposes into direct sum of irreps as follows:
@@
This induces a block structure in $\hat{{\bm{W}}}_{l}$, where the $(\psi,i;\psi^{\prime},j)$-th block block maps the $i$-th input block transforming according to $\psi$ to the $j$-th output block transforming under $\psi^{\prime}$, with $i\in[m_{l-1,\psi}]$ and $j\in[m_{l,\psi^{\prime}}]$. By Theorem A.6, there are no linear maps equivariant to $\psi$ and $\psi^{\prime}$ whenever these representations are in-equivalent; this implies the matrix $\hat{{\bm{W}}}_{l}$ is sparse since its $(\psi,i;\psi^{\prime},j)$-th block is non-zero only when $\psi\cong\psi^{\prime}$. Then, the non-zero $(\psi,i,j)$-th block (here, we drop the redundant $\psi^{\prime}$ index) is denoted by $\hat{w}_{i,j}(\psi)$ and should commute with $\psi(g)$ for any $g\in G$. This is achieved by expressing this matrix as a linear combination of a few fixed basis matrices spanning the space; see [behboodi_pac-bayesian_2022, cesa2022a] for more details.
@@
Each module in an equivariant network should commute with the action of the group on its own input in order to guarantee the overall equivariance of the model. Hence, the activation function $\sigma$ used in the intermediate layer should be equivariant with respect to $\rho_{1}(g)$ as well, i.e., $\sigma(\rho_{1}(g){\bm{x}})=\rho_{2}(g)\sigma({\bm{x}}),\forall g\in G$. Different activation functions are used in the literature. As we use unitary representations, one can use norm nonlinearities, $\sigma({\bm{x}})=\eta(\left\|{\bm{x}}\right\|)\frac{{\bm{x}}}{\left\|{\bm{x}}\right\|}$ for a suitable function $\eta:\mathbb{R}^{+}_{0}\to\mathbb{R}^{+}_{0}$. It can be seen that $\sigma(\rho_{1}(g){\bm{x}})=\rho_{1}(g)\sigma({\bm{x}})$. Some examples are norm ReLU [Worrall2017-HNET], squashing [sabour_dynamic_2017] and gated nonlinearities [Weiler2018-STEERABLE].
@@
The last layer maps the features after the activation to a single scalar and is assumed to be invariant. This linear layer can be thought of as a special case of the general framework above, with the choice of output representation $\rho_{2}$ being a collection of trivial representations, i.e., $\rho_{2}(g)=I$ is the identity. Since the trivial representation is an irrep, by Theorem A.6, this linear map only captures the invariant information in the output of the activation function. In the case of a GCNN with pointwise activation $\sigma$, this layer acts as an averaging pooling operator over the $|G|$ entries of each channel and learns the weights to linearly combine the $c_{1}$ independent channels. Note that one can also include a custom pooling operator $P(\cdot)$ (e.g., max-pooling) between the activation layer $\sigma$ and the final layer ${\bm{u}}$ like in eq. 3: this gives additional freedom to construct the final invariant features beyond just average pooling.